1.     In a survey of 703 randomly selected workers, 61% got their jobs through networking (based on data from Taylor Nelson Sofres Research).  Use the sample data with a 0.05 significance level to test the claim that most (more than 50%) workers get their jobs through networking.  What does the result suggest about the strategy for finding a job after graduation?



1
Expert's answer
2021-11-17T06:21:18-0500

Ho: p=0.5

Ha: p>0.5

The test statistic is

"Z=(phat-p)/ \sqrt{(p*(1-p)}/n)\\

=(0.61-0.5)/ \sqrt{(0.5*0.5}/703)\\

=5.83"

Given a=0.05, the critical value is Z(0.05)=1.645 (from standard normal table)

Since Z=5.83 is larger than 1.645, we reject Ho.

So we can conclude that most workers get their jobs through networking.